OpenStudy (anonymous):

Juice has 30% oj and another has 10% oj. How many liters of each is needed to get 20L at 16% oj?

OpenStudy (anonymous):

posted above my plea

OpenStudy (mertsj):

x= amount of 30% 20-x= amount of 10% .3x+.1(20-x)=.16(20)

OpenStudy (mertsj):

Do you understand the equation and can you solve it?

OpenStudy (anonymous):

kind of understand - getting confused a little

OpenStudy (anonymous):

greatly due to frustration I am sure

OpenStudy (mertsj):

If you have 20 ounces of candy and 50% of it is M&M's then .5(20) = amount of M&M's so if you have x L of solution and it is 30% OJ, then .3(x)= amount of OJ

OpenStudy (mertsj):

Starting amount of OJ + added amount OJ= final amount of OJ

OpenStudy (mertsj):

\[.3x+.1(20-x)=.16(20)\] \[.3x+2-.1x=3.2\] \[.2x=1.2\] x=6 L of 30 % OJ 20-6=14 L of 10% OJ

OpenStudy (anonymous):

plz dont laugh - on the second line where did the 2 come from - lookin at my work I think that is where I was messing up

OpenStudy (phi):

Juice has 30% OJ and another has 10% OJ. How many liters of each is needed to get 20L at 16% OJ? To answer these “mixture problems”, we need to set up some equations. Once we have the equations, we solve them. Part I What are the equations? Start with the question “How many liters of each is needed to get 20L?” Think, “I add some of juice A and some of juice B to get 20 L” which is written as A+B= 20 We don’t know the amounts (yet), but it should make sense. A liters from the 30% juice plus B liters from the 10% juice must equal 20 liters. Next, the key idea or insight. If you have 30% OJ this means 30% of the juice is pure OJ and 70% (100% - 30%) is pure water. Use this idea to figure, if I have 20 liters of 16% OJ, then 16% of 20 liters is pure OJ. As an equation: 0.16*20 liters= 3.2 liters pure OJ Apply this same idea to A liters of the 30% OJ: if I have A liters of 30% juice, then I have 0.30*A of pure OJ. And for B liters of the 10% OJ, I have 0.1B liters of pure OJ. So, thinking just about the (pure) OJ: 0.3A + 0.1B = 3.2 This says that the amount of OJ in A liters of 30% juice plus the amount of OJ in B liters of 10% juice must add up to 3.2 liters in the final mix. Part II Solve the equations 0.3A+0.1B= 3.2 A+B= 20 multiply the first equation by 10, to get rid of the decimals. We get 3A+B= 32 A +B= 20 Subtract the two equations, term by term (3A-A)+(B-B)=(32-20) 2A= 12 A= 6 Now use one of the two equations to find B. Use the simpler one, A+B=20 6+B=20 B=14 The final answer is, 6 liters of the 30% juice, and 14 liters of the 10% juice.